# CAMEO: Collection of Multilingual Emotional Speech Corpora

### v2

@@ 5 Benchmark Results
Table 6 summarizes the overall results, which were measured using the macro-averaged F1 score, weighted F1 score, and accuracy. As expected, the overall performance was limited. The best results were obtained with Qwen2-Audio, which achieved an F1 macro score of only 0.2 for temperature of 0. This further indicates the difficulty of this task in a zero-shot approach.
@@
Table 4 shows the F1 macro scores obtained by the models at a given temperature setting for each language. Qwen2-Audio consistently outperformed the other systems across all languages. The notably higher scores achieved by Qwen2-Audio on English, French, and German and by Ultravox on French and German raise the question of possible contamination.
This concern was investigated further, and F1 macro scores were calculated for the datasets. The analysis revealed that Qwen2-Audio performed well on CREMA-D (0.80), eNTERFACE (0.54), and RAVDESS (0.73). These results suggest that these datasets were encountered during training or fine-tuning. Similarly, Ultravox exhibited superior scores on both eNTERFACE (0.79) and Oréau (0.58), which may imply a similar overlap.
Table 5 shows the F1 scores across seven primary emotional states. Interestingly, the systems performed best for sadness, and worst for neutrality, despite the narrative nature of neutral speech.

### v3

@@ 5 Benchmark Results
Table 6 reports the overall performance using macro-averaged F1 score, weighted F1 score, and accuracy. Performance remained low across all models and temperatures, highlighting the difficulty of multilingual speech emotion recognition in a zero-shot setting. The best-performing system was Qwen2-Audio, which achieved a macro F1 score of 0.20 at a temperature of 0.0. Though modest, this result is consistent with the absence of task-specific fine-tuning and emphasizes the difficulty of mapping acoustic emotion cues to specific labels across languages using general-purpose AudioLLMs.
@@
Table 4 shows the macro F1 scores for each language at a fixed temperature. Qwen2-Audio consistently outperformed the other models across all evaluated languages, suggesting stronger generalization capabilities or broader exposure to emotionally annotated speech during pretraining. Qwen2-Audio demonstrated notably higher performance for English, French, and German. Ultravox also demonstrated higher performance for French and German, albeit to a lesser extent. These results suggest the possibility of data contamination, especially since these languages are well represented in publicly available emotional speech corpora.
To investigate potential overlap effects, macro F1 scores were computed at the dataset level. Qwen2-Audio performed substantially better on CREMA-D (0.80), RAVDESS (0.73), and eNTERFACE (0.54). A similar pattern was observed for Ultravox, which showed elevated performance on eNTERFACE (0.79) and Oréau (0.58). Since these datasets are widely used benchmarks in SER, the results suggest that they were likely encountered during training or fine-tuning.
Table 5 shows F1 scores for the sever primary emotional states. Sadness was the most reliably recognized emotion across all systems, while neutral speech was the most challenging. This may be due to the fact that sadness often exhibits more stable and distinctive acoustic cues, such as reduced pitch variation and slower tempo. In contrast, neutral speech is inherently heterogeneous and often acoustically overlaps with low-arousal emotional states. Despite its narrative or conversational nature, the lack of salient prosodic markers likely contributes to frequent misclassification.